OpenStudy (anonymous):

find the limit. lim(θ→0)(sin3x/(5x^3-4x)

OpenStudy (anonymous):

Do you mean\[\lim_{\theta\to0}\frac{\sin3\theta}{5\theta^3-4\theta}\]

OpenStudy (dan815):

There is no theta

OpenStudy (anonymous):

yes i do mean theat

OpenStudy (anonymous):

*theta

OpenStudy (anonymous):

\[\begin{align*}\lim_{\theta\to0}\frac{\sin3\theta}{5\theta^3-4\theta}&=\lim_{\theta\to0}\frac{\sin3\theta}{\theta(5\theta^2-4)}\\ &=\lim_{\theta\to0}\frac{3\sin3\theta}{3\theta(5\theta^2-4)}\\ &=\lim_{\theta\to0}\frac{\sin3\theta}{3\theta}\cdot\lim_{\theta\to0}\frac{3}{5\theta^2-4} \end{align*}\]